Three times the sum of three consecutive intergers is equal to the produsct of the two smallest integers.
Answer by richwmiller(17219)

3*(n+n+1+n+2)=n*(n+1)
3*(3n+3)=n^2+n
9n+9=n^2+n
n=-1 and n=9
-1,0,1
check
0=0
ok
9,10,11
check
3*(30)=9*10
90=90
so there are two sets of answers
-1,0,1
and 9,10,11